OpenStudy (anonymous):

SAVE MEH 1. x^-1y^-2 over z^-3 (idk how to make a fraction) 2. 2a^4 over 5b^-2 3. a^6n over a^3n 4. (5a^3n)^3 5. (x^n+5 over y) ^2 6. find the value when x=2 and y=3 a. 2x^0y^-2; (2x+5y)^0; and x^-3y^-3 that's it for today I promiseee

OpenStudy (anonymous):

6. \[\Large 2x^0y^{-2} \longrightarrow 2\cdot 1\cdot 3^{-2}\longrightarrow 2\cdot \frac{1}{3^2} ...\]

OpenStudy (anonymous):

the important thing here is: \[\frac{1}{x^a} = x^{-a}\] for the first one: 1. x^-1y^-2 over z^-3 \[\frac{x^{-1}y^{-2}}{z^{-3}} = \frac{1}{x}\frac{y^{-2}}{z^{-3}} = \frac{1}{x}\frac{1}{y^2}\frac{1}{z^{-3}}=\frac{1}{x}\frac{1}{y^2}\frac{z^3}{1}\]\[=\frac{z^3}{xy^2}\]

OpenStudy (anonymous):

6. \[\Large (2x+5y)^0=\text{Any number \in power 0 is equal \to 1} \] \[\LARGE (2x+5y)^0=1\]\ because... \[\LARGE a^0=a^{k-k}=\frac{a^k}{a^k}=1\]

OpenStudy (anonymous):

3. \[\LARGE \frac{a^{6n}}{a^{3n}}=a^{6n-3n}=a^{3n}\]

OpenStudy (anonymous):

thank you for actually answering them kreshnik

OpenStudy (anonymous):

hold on... I'll do more..

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

2. \[\LARGE \frac{2a^4}{5b^{-2}}\] this one doesn't simplifieis any further !! only if you want to write it like...\\\[\LARGE \frac{2a^4\cdot b^{2}}{5}\] 5. \[\LARGE \left(\frac{x^n+5}{y}\right)^2=\frac{(x^n)^2+2\cdot 5x^n+5^2}{y^2}=\frac{x^{2x}+10x^n+25}{y^2}\]

OpenStudy (anonymous):

I have to go right now but I WILL look at hese later THANK YOU SOOO MUCH

OpenStudy (anonymous):

=(x^(2x)+10x^n+25)/(y^2) --------- this is the answer... maybe you don't see it... I'm going now bye ;)

OpenStudy (anonymous):

which one you need to do now? ..where are you stuck

OpenStudy (lgbasallote):

lol...@rebeccaskell94 is so loved by OS =))))

OpenStudy (anonymous):

try something at least ! show us what you've done so far! @rebeccaskell94

OpenStudy (callisto):

Okay.. the first time when i nearly finished typing, my dearest google just said sth went wrong, and all my things are gone! Start again... 1. x^-1y^-2 over z^-3 = z^3 / xy^2 2. 2a^4 over 5b^-2 = 2a^4b^2 /5 3. a^6n over a^3n = a^(6n-3n) = a^ (3n) 4. (5a^3n)^3 = 5^3[a^(3n)]^3 = 125a^9n 5. (x^n+5 over y) ^2 = (x^n +5)^2 / y^2 = (x^2n + 10x^n +25) / y^2 6. find the value when x=2 and y=3 a. 2x^0y^-2; (2x+5y)^0; and x^-3y^-3 Put x=2, and y=3 into a. 2x^0y^-2 ; (2x+5y)^0 x^-3y^-3 = 2(2^0)(3^-2) ; (2(2)+5(3))^0 =2^-3 (3^-3) Can you solve 6 your own? if you understand 1-5, you can do 6 ...i think..